In a Young’s double slit interference experiment, the slits are at a distance 2L from each
other and the screen is at a distance D from the slits. If a glass slab of refractive index 
and thickness d is placed in the path of one of the beams, the minimum value of d for
the central fringe to be dark is
1
Expert's answer
2020-10-13T18:06:15-0400

Youngs double slit experiment

"\Delta x =S"2"*P-S" 1*P

from pathegorian relationship"\Delta x=d \sin \theta"

"\sin \theta" ="\dfrac{Xn}{S}"

"\therefore" "\Delta x =" "\dfrac{Xnd}{S}"

from constructive interference "\Delta x =m \lambda"

"\therefore m \lambda =" "\dfrac{Xn d}{S}"

xn ="\dfrac{m \lambda}{d}"